OpenStudy (anonymous):

jill the animal trainer keeps only two types of animals, horses which have four feet and ostriches which have two feet. jill who is always looking for fun math problems notices that in her corral she can count 25 heads and 72 legs.. how many ostriches and horses does she have?? describe the solution to this problem using two diffrent problem solving strategies.

OpenStudy (anonymous):

There are 25 heads, so the total number of animals must be 25. Let the number of horses be x. So, number of ostriches = 25 - x 4x + 2(25 - x) = 72 4x + 50 - 2x = 72 2x + 50 = 72 2x = 22 x = 11 25 - x = 14

OpenStudy (anonymous):

Am I right?

OpenStudy (anonymous):

is this the only way to solve this?

OpenStudy (anonymous):

I think so. You have only one equation and two variables. So, there is no other choice than to substitute one of the variables in terms of the other one.